# Numerical solution of elliptic distributed optimal control problems with boundary value tracking

### v1

@@ 1 Introduction
where $\overline{y}\in L^{2}(\Gamma)$ denotes a given target, $\varrho\in{\mathbb{R}}_{+}$ is a positive regularization or cost parameter, and $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a bounded Lipschitz domain with the boundary $\Gamma=\partial\Omega$. Our work was inspired by the paper [4], where the $L^{2}$ regularization corresponding to the choice $U=L^{2}(\Omega)$ was investigated. The boundary tracking is a special case of the partial tracking of a given target in a subset of $\Omega$ that is also called limited observation; see, e.g., [5]. In this contribution, we consider the energy regularization corresponding to the choice $U=\widetilde{H}^{-1}(\Omega):=[H^{1}(\Omega)]^{*}$, where we use the standard notations for Lebesgue and Sobolev spaces; see, e.g., [8]. In order to follow the abstract theory presented in [3], we define the state space in such a way that the state-to-control map is an isomorphism. This allows us to derive a state-based formulation which is the basis for the numerical solution. Here we restrict the analysis to a conforming tensor-product finite element (fe) discretization that finally leads to a linear system of algebraic equations for which fast solvers can be constructed.
@@ 2 State-based variational reformulation
The variational formulation of the Neumann bvp (2) reads to find $y_{\varrho}\in H^{1}(\Omega)$ such that
@@
is satisfied for all $y\in H^{1}(\Omega)$, where we assume $u_{\varrho}\in\widetilde{H}^{-1}(\Omega)$. While the Neumann boundary condition in (2) enters the variational formulation (3) in a natural way, this condition has to be included in the definition of the state space
@@
Note that the extension $\overline{y}_{e}\in Y$ of $\overline{y}\in H^{1/2}(\Gamma)$ as used in Lemma 2.1 can be computed as solution of the first boundary value problem for the biharmonic equation,
$\Delta^{2}\overline{y}_{e}=0\quad\mbox{in}\;\Omega,\quad\overline{y}_{e}=\overline{y},\quad\partial_{n}\overline{y}_{e}=0\quad\mbox{on}\;\Gamma.$ (9)
@@ 3 Conforming FE discretization on tensor product meshes
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)},$ (12)
while, for $\overline{y}\in H^{1/2}(\Gamma)$ and choosing $\varrho=h$, we have
@@
A related result holds for $\overline{y}\in H^{1}(\Gamma)$ where, for $\varrho=h$,
@@
Next we consider the case when $\overline{y}_{e}\in Y$ is the extension of $\overline{y}$ as solution of the biharmonic boundary value problem (9). When using the triangle inequality twice in combination with Cea’s estimate (11), we can write, for arbitrary $y_{h}\in Y_{h}$
@@
In particular for $y_{h}=P_{h}\overline{y}_{e}\in Y_{h}$ being the $L^{2}$ projection of $\overline{y}_{e}$, we have the standard fe error estimates
@@
Finally, we consider the case $\overline{y}\in H^{1}(\Gamma)$ which implies $\overline{y}_{e}\in H^{3/2}(\Omega)$. In this case, we have
@@
and
$\|\overline{y}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}.$
Together with (8) we therefore conclude
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho^{2}+\varrho h+h^{2})\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}^{2},$
and for $\varrho=h$ this gives (14). ∎
@@
Since $\overline{y}_{e}$ is defined as solution of the biharmonic boundary value problem (9), we have to formulate additional compatibility conditions in order to ensure $\overline{y}_{e}\in H^{2}(\Omega)$, when $\Gamma=\partial\Omega$ is piecewise smooth. In particular, the tangential derivatives of $\overline{y}$ have to be in $H^{1/2}(\Gamma)$, see, e.g., the discussion in [2, 6] in the two-dimensional case. Then we can use (3) to conclude
@@ 4 Fast solvers
Eliminating $\mathbf{y}_{I}=-\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}\mathbf{y}_{B}$ from the linear system (19), we arrive at the boundary Schur complement system
@@
with $h$ and $\varrho$ independent positive constants $\widetilde{c}$ and $\mathring{c}$. The choice $\varrho\leq h$ delivers the desired result. It is recommended to use $\mbox{lump}(\mathbf{M}_{BB})$ as diagonal preconditioner in the Preconditioned Conjugate Gradient (PCG) method since it provides the right scaling. The numerical results presented in Section 5 show that the system (19) can also efficiently be solved by means of PCG with a simple Algebraic MultiGrid (AMG) preconditioner.
@@ 5 Numerical results
on the boundary $\Gamma=\partial\Omega$ of the domain $\Omega=(0,1)^{3}$. We mention that $\overline{y}$ is the trace of a smooth function with vanishing normal derivative on the boundary $\Gamma$, i.e., we have $\overline{y}\in H^{2}(\Gamma)$, and the error estimate (18) applies when choosing $\varrho=h^{2}$. We use a tensor product mesh as described in Section 3. The initial mesh contains $5$ vertices in each direction, and $125$ in total with mesh size $h=0.25$. We note that we have only $3$ dofs in each direction, and $27$ in total for the initial level. Table 1 provides the numerical results starting from level $\ell=1$ with $27$ dofs and running to the finest discretization level $\ell=7$ obtained by $6$ uniform refinements of the initial mesh. The fourth column displays the $L^{2}$ error $\|y-\overline{y}\|_{L^{2}(\Gamma)}$ on the boundary. As expected, we observe second order of convergence; cf. experimental order of convergence (eoc) given in the fifth column. We first solve the original system (19) by means of AMG preconditioned CG iterations (#AMG-PCG its), and observe that not more than $4$ iterations are needed in order to reach a relative residual error of $10^{-9}$. We further test the CG and lumped mass preconditioned CG solvers for the Schur complement equation (20) until the relative residual error reaches $10^{-9}$. The number of Schur complement CG (#SCG its) and lumped mass preconditioned CG (#SPCG its) iterations are displayed in the last two columns of Table 1. As expected from the theoretical results given in Section 4, we see level-independent iteration numbers in both cases. Moreover, the lumped-mass preconditioner further reduces the number of iteration by the scaling effect. We note that the action of $\mathring{\mathbf{K}}_{II}^{-1}$ to a vector within the multiplication of the Schur complement $\mathbf{S}_{BB}$ by some vector (iterate) is realized by an AMG preconditioned CG method until the relative residual error is reduced by a factor $10^{10}$. The latter accuracy of this inner PCG iteration can be adapted (reduced !) to the outer CG/PCG iteration following the results from [7].
@@
Table 1: Target (22): error = $\|y-\overline{y}\|_{L^{2}(\Gamma)}$, number of AMG CG iterations (#ACG-PCG its) for the original system (19), and number of CG (#SCG its) and lumped-mass preconditioned CG (#PSCG its) iterations for the Schur complement system (20), $\varrho=h^{2}$.
@@
is a trace of a smooth function which does not fulfill the homogeneous Neumann boundary conditions on the boundary $\Gamma$ of $\Omega=(0,1)^{3}$, i.e., $\overline{y}\in H^{3/2-\varepsilon}(\Gamma)$, and $\overline{y}_{e}\in H^{2-\varepsilon}(\Omega)$, $\varepsilon>0$. Similar as in (17) we therefore expect a reduced eco of about 1.5. We perform the same tests for the target (23) as for the previous example for the target (22). The results are given in Table 2, where we have used $\varrho=h^{3/2}$ as prescribed by the theory. The iteration numbers are again independent of $h$, and show the same behavior as in the case of the first example.
@@ Acknowledgement
Let $\Gamma=\partial\Omega$ for $\Omega=(0,1)^{d}$. Assume $y\in H^{2}(\Gamma)$, i.e., there exists an extension $y_{e}\in Y\cap H^{5/2}(\Omega)$ as solution of the biharmonic boundary value problem (9). Then, there holds the error estimate

### v2

@@ 1 Introduction
where $\overline{y}\in L^{2}(\Gamma)$ denotes a given target, $\varrho\in{\mathbb{R}}_{+}$ is a positive regularization or cost parameter, and $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a bounded Lipschitz domain with the boundary $\Gamma=\partial\Omega$. Our work was inspired by the paper [4], where the $L^{2}$ regularization corresponding to the choice $U=L^{2}(\Omega)$ was investigated. The boundary tracking is a special case of the partial tracking of a given target in a subset of $\Omega$ that is also called limited observation; see, e.g., [5]. There are many important practical applications, e.g., inverse heat transfer, photoacoustic tomography, etc., see also [4]. In this contribution, we consider the energy regularization corresponding to the choice $U=\widetilde{H}^{-1}(\Omega):=[H^{1}(\Omega)]^{*}$ such that the solution of (2) defines an isomorphism when considering $y_{\varrho}\in H^{1}(\Omega)$. When using $U=L^{2}(\Omega)$ this implies $y_{\varrho}\in H^{1}_{\Delta}(\Omega):=\{y\in H^{1}(\Omega):\Delta y\in L^{2}(\Omega)\}$ to ensure an isomorphism, which afterwards requires higher order basis functions for a conforming discretization. Here we use the standard notations for Lebesgue and Sobolev spaces; see, e.g., [8]. In order to follow the abstract theory presented in [3], we define the state space in such a way that the state-to-control map is an isomorphism. This allows us to derive a state-based formulation which is the basis for the numerical solution. Here we restrict the analysis to a conforming tensor-product finite element (fe) discretization that finally leads to a linear system of algebraic equations for which fast solvers can be constructed. Note that $c$ denotes a universal constant independent of the regularization parameter $\varrho$, and the discretization parameter $h$.
@@ 2 State-based variational reformulation
The variational formulation of the Neumann boundary value problem (2) reads to find $y_{\varrho}\in H^{1}(\Omega)$ such that
@@
is satisfied for all $y\in H^{1}(\Omega)$, where we assume $u_{\varrho}\in\widetilde{H}^{-1}(\Omega)$. The variational problem has a unique solution due to Lax-Milgram’s lemma; see, e.g., [8]. While the Neumann boundary condition in (2) enters the variational formulation (3) in a natural way, this condition has to be included in the definition of the state space
@@
Here, the constant $c$ depends on the mapping properties of the partial differential operator $I-\Delta$ in certain Sobolev spaces, and the trace theorem.
@@ 3 Conforming FE discretization on tensor product meshes
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)}.$ (11)
If $\overline{y}\in H^{1/2}(\Gamma)$ is the Dirichlet trace of $\overline{y}_{e}\in Y$, and choosing $\varrho=h$, we have
@@
For $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1}(\Gamma)$, $\overline{y}_{e}\in H^{3/2+\varepsilon}(\Omega)$ with $\varepsilon>0$, and choosing $\varrho=h$, we have
@@
For $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1/2}(\Gamma)$ we use the triangle inequality twice and Cea’s estimate (10) for arbitrary $y_{h}\in Y_{h}$ to write
@@
In particular, for $y_{h}=P_{h}\overline{y}_{e}\in Y_{h}$ being the $L^{2}$ projection of $\overline{y}_{e}$, we have the standard fe error estimates; see, e.g., [8],
@@
Finally, we consider the case $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1}(\Gamma)$ with $\overline{y}_{e}\in H^{3/2+\varepsilon}(\Omega)$ for some $\varepsilon>0$. In this case, we have
@@
and   $\|\overline{y}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}.$ Together with (8) we therefore conclude $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho^{2}+\varrho h+h^{2})\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}^{2},$ and, for $\varrho=h$, this gives (13). ∎
@@
Note that $\overline{y}_{e}\in H^{2}(\Omega)$ implies some additional compatibility conditions on $\overline{y}$, when $\Gamma=\partial\Omega$ is piecewise smooth. In particular, the tangential derivatives of $\overline{y}$ have to be in $H^{1/2}(\Gamma)$, see, e.g., the discussion in [2, 6] in the two-dimensional case. Then we can use (3) to conclude
@@ 4 Fast solvers
Eliminating $\mathbf{y}_{I}=-\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}\mathbf{y}_{B}$ from the linear system (18), we arrive at the boundary Schur complement (SC) system
@@
with $h$ and $\varrho$ independent positive constants $\widetilde{c}$ and $\mathring{c}$ arising from the estimates $\lambda_{max}(\mathbf{M}_{BB}^{-1}\widetilde{\mathbf{K}}_{BB})\leq\widetilde{c}h^{-1}$ and $\lambda_{max}(\mathbf{M}_{BB}^{-1}\mathring{\mathbf{S}}_{BB})\leq\mathring{c}h^{-1}$ of the maximal eigenvalues of $\mathbf{M}_{BB}^{-1}\widetilde{\mathbf{K}}_{BB}$ and $\mathbf{M}_{BB}^{-1}\mathring{\mathbf{S}}_{BB}$, respectively. The choice $\varrho\leq h$ delivers the desired result. It is recommended to use $\mbox{lump}(\mathbf{M}_{BB})$ as diagonal preconditioner in the Preconditioned Conjugate Gradient (PCG) method since it provides the right scaling. The numerical results presented in Section 5 show that the system (18) can also efficiently be solved by means of PCG with a simple Algebraic MultiGrid (AMG) preconditioner.
@@ 5 Numerical results
on the boundary $\Gamma=\partial\Omega$ of the domain $\Omega=(0,1)^{3}$. We mention that $\overline{y}$ is the trace of a smooth function with vanishing normal derivative on the boundary $\Gamma$, i.e., we have $\overline{y}\in H^{2}(\Gamma)$, and the error estimate (17) applies when choosing $\varrho=h^{2}$. We use a tensor product mesh as described in Section 3. The initial mesh contains $5$ vertices in each direction, and $125$ in total with mesh size $h=0.25$. We note that we have only $3$ dofs in each direction, and $27$ in total for the initial level. Table 1 provides the numerical results starting from level $\ell=1$ with $27$ dofs and running to the finest discretization level $\ell=7$ obtained by $6$ uniform refinements of the initial mesh. The fourth column displays the $L^{2}$ error $\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}$ between the computed fe solution $y_{\ell}=y_{\varrho_{\ell}h_{\ell}}$ and the target $\overline{y}$ on the boundary. As expected, we observe second order of convergence; cf. experimental order of convergence (eoc) given in the fifth column. We first solve the original system (18) by means of AMG preconditioned CG iterations (#AMG-PCG its), and observe that not more than $4$ iterations are needed in order to reach a relative residual error of $10^{-9}$. We further test the CG and lumped mass preconditioned CG solvers for the Schur complement equation (19) until the relative residual error reaches $10^{-9}$. The number of Schur complement CG (#SCG its) and lumped mass preconditioned CG (#SPCG its) iterations are displayed in the last two columns of Table 1. As expected from the theoretical results given in Section 4, we see level-independent iteration numbers in both cases. Moreover, the lumped-mass preconditioner further reduces the number of iteration by the scaling effect. We note that the action of $\mathring{\mathbf{K}}_{II}^{-1}$ to a vector within the multiplication of the Schur complement $\mathbf{S}_{BB}$ by some vector (iterate) is realized by an AMG preconditioned CG method until the relative residual error is reduced by a factor $10^{10}$. The latter accuracy of this inner PCG iteration can be adapted (reduced !) to the outer CG/PCG iteration following the results from [7].
@@
Table 1: Target (21): error = $\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}$, eoc = $\log_{2}(\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}/\|y_{\ell+1}-\overline{y}\|_{L^{2}(\Gamma)})$, number of AMG CG iterations (#AMG-PCG its) for the original system (18), and number of CG (#SCG its) and lumped-mass preconditioned CG (#PSCG its) iterations for the SC system (19), $\varrho=h^{2}$.
@@
is a trace of a smooth function which does not fulfill the homogeneous Neumann boundary conditions on the boundary $\Gamma$ of $\Omega=(0,1)^{3}$, i.e., $\overline{y}\in H^{3/2-\varepsilon}(\Gamma)$, and $\overline{y}_{e}\in H^{2-\varepsilon}(\Omega)$, $\varepsilon>0$. Similar as in (16) we therefore expect a reduced eoc of about 1.5. We perform the same tests for the target (22) as for the previous example for the target (21). The results are given in Table 2, where we have used $\varrho=h^{3/2}$ as prescribed by the theory. The iteration numbers are again independent of $h$, and show the same behavior as in the case of the first example.
@@
Finally, we consider a discontinuous target $\overline{y}$ that is equal to 1 inside the region $0.25,0.75]\times[0.25,0.75]$ and 0 outside on each face $\Gamma_{i}\subset\Gamma=\partial\Omega,i=1,\dots,6$ the unit cube $\Omega=(0,1)^{3}$. We perform the same tests as for the previous examples. The results are given in Table 3, where we have used $\varrho=h$ as prescribed by the theory. We observe the expected convergence rate eoc = 0.50. Furthermore, the iteration numbers are again independent of $h$, and show the same behavior as in the case of the preceding two examples. Figure 2 shows the computed state $y_{\ell}$ on the surface $\Gamma$ of the cube $\Omega$ (left) and on three cuts through the cube.
$\ell$ #Dofs $h$ error eoc #AMG-PCG its #SCG its #PSCG its
11 $27$ $2^{-2}$ 9.8179.817e2-2 - 33 33 33
22 $343$ $2^{-3}$ 6.5356.535e1-1 - 44 $10$ $10$
33 $3,375$ $2^{-4}$ 4.7684.768e1-1 $0.45$ 55 $17$ $17$
44 $29,791$ $2^{-5}$ 3.5303.530e1-1 $0.43$ 44 $22$ 1919
55 $250,047$ $2^{-6}$ 2.5402.540e1-1 $0.47$ 44 $23$ 2121
66 $2,048,383$ $2^{-7}$ 1.8091.809e1-1 $0.49$ 44 $23$ 2121
77 $16,581,375$ $2^{-8}$ 1.2831.283e1-1 $0.50$ 55 $23$ 2121
Table 3: Discontinuous target with 1 inside the region of $0.25\quad 0.75]\times[0.25\quad 0.75]$ and 0 outside this region on each face of the unique cube: Same agenda as Table 1, but now $\varrho=h$.
Figure 2: Computed state for the discontinuous target at level 77.
@@ Acknowledgement
Let $\Gamma=\partial\Omega$ for $\Omega=(0,1)^{d}$. Assume $y\in H^{2}(\Gamma)$, i.e., there exists an extension $y_{e}\in Y\cap H^{5/2}(\Omega)$. Then, there holds the error estimate